# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v3

@@ Abstract
Missing data is a pervasive problem in tabular settings. Existing solutions range from simple averaging to complex generative adversarial networks, but due to each method’s large variance in performance across real-world domains and time-consuming hyperparameter tuning, no default imputation method exists. Building on TabPFN, a recent tabular foundation model for supervised learning, we propose TabImpute, a pre-trained transformer that delivers accurate and fast zero-shot imputations requiring no fitting or hyperparameter tuning at inference-time. To train and evaluate TabImpute, we introduce (i) an entry-wise featurization for tabular settings, which enables a $100\times$ speedup over the previous TabPFN imputation method, (ii) a synthetic training data generation pipeline incorporating realistic missingness patterns, and (iii) MissBench, a comprehensive benchmark with $42$ OpenML datasets and $13$ new missingness patterns. MissBench spans domains such as medicine, finance, and engineering, showcasing TabImpute’s robust performance compared to $12$ established imputation methods.
@@ 1 Introduction
Missing data is ubiquitous across tabular datasets, affecting statisticians, economists, health officials, and businesses. For example, healthcare datasets may lack some recorded blood pressure measurements, or datasets merged from multiple sources may only share partial features. Regardless of the source, missing data must be imputed to numerical values before employing statistical or machine learning models. Imputation methods range from simple approaches (e.g., constant values and averages) to more sophisticated techniques like nearest-neighbor-based methods (Batista01052003), matrix factorization approaches such as SoftImpute (hastie2015matrix), and random forest regression, notably the MissForest algorithm (missforest2011). However, each method in the literature is typically tailored for specific settings, with performance varying significantly across datasets, domains, and missingness patterns (van2012flexible; jarrett2022hyperimpute; agarwal2023causal; Ibrahim01032005). Building on recent advances in tabular representation learning (hollmann2023tabpfn; ye2025closer), we propose a pre-trained transformer specifically designed for the tabular missing-data problem that produces accurate and fast zero-shot imputations.
@@
Figure 1: Evaluation on real-world OpenML data: MissBench. We compare TabImpute with $12$ other popular methods on MissBench. In panel (a), we plot the imputation accuracy (defined as 1 - normalized RMSE), which is calculated for each method, normalized within a dataset, and averaged across datasets and $13$ missingness patterns. Error bars indicate 95% confidence intervals. In panel (b), we compare the runtime per table entry. Any method not labeled (GPU) is tested on a CPU because that method is not implemented for GPUs. See Sec. 3 for our exact computing specifications and Sec. 4 for accuracy score methodology.
@@
rubin1976inference proposed analyzing missingness based on its relationship with the data-generating process to determine whether missingness biases downstream estimation. Rubin demonstrated that when missingness operates independently of the underlying data, the observed data distribution provides an unbiased foundation for estimation. This framework categorizes missingness into three classes: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR) (van2012flexible; sportisse2020aimputation; sportisse2020b). MCAR defines scenarios where missingness occurs uniformly and independently of all data values. MAR encompasses cases where missingness depends on observed variables that can be appropriately conditioned upon during analysis. MNAR describes situations where missingness depends on unobserved factors that cannot be easily conditioned on.
@@
While this framework characterizes how missingness relates to the underlying data-generating process, it largely treats the missingness indicators for each variable as conditionally independent across entries. Recent work on structured missingness argues that, especially in large-scale multi-source datasets, missingness itself can exhibit rich multivariate structure that is not captured by the standard MCAR/MAR/MNAR taxonomy. In particular, mitra2023learning introduce structured missingness as an umbrella term for mechanisms in which missingness follows systematic patterns. jackson2023completecharacterisationstructuredmissingness provide a complementary characterization that allows dependencies among missingness indicators for different variables.
@@
Previous work typically introduced new missingness patterns within these categories and proposed pattern-specific solutions. For instance, agarwal2023causal proposed a nearest neighbor-based matrix completion method specifically designed for a block-wise MNAR pattern. We instead develop a single method that performs well across diverse patterns and data domains by building on TabPFN, a popular tabular foundation model for supervised learning.
@@
TabPFN is a pre-trained transformer model for supervised learning that performs well across a variety of domains without any fine-tuning (hollmann2025accurate). The team behind TabPFN created an imputation method in their tabpfn-extensions Python package by using the TabPFN model in iterative column-wise imputation (available on GitHub [^1] https://github.com/PriorLabs/tabpfn-extensions/blob/main/src/tabpfn_extensions/unsupervised/unsupervised.py). However, when evaluated on our new benchmark MissBench–consisting of $42$ real-world OpenML datasets and $13$ missingness patterns–TabPFN’s imputation approach shows low accuracy and slow runtime. We improve on this approach by introducing a new entry-wise featurization (EWF), allowing parallel prediction of each missing value using TabPFN’s model, denoted EWF-TabPFN. To further improve on EWF-TabPFN, we train a new underlying model specifically designed for tabular data imputation, TabImpute, to better fit this class of tasks, achieving the imputation accuracy shown in Fig. 1 and Tab. 1.
@@
  • We propose a new state-of-the-art pre-trained transformer model, TabImpute, for missing data imputation by building on TabPFN and introducing a new entry-wise missing data featurization (see Sec. 3.1 for details).
  • We develop a synthetic data generation pipeline to create training datasets with a comprehensive collection of missing values covering a wide range of MNAR patterns (see Sec. 3.2, Sec. A.2, and Tab. 10 for details).
  • We demonstrate TabImpute’s performance on a novel, comprehensive test bench, denoted MissBench, using $42$ real-world OpenML datasets and $13$ missingness patterns (see Sec. 4 for details). Our benchmark builds on previous work by testing more datasets and more missingness patterns: HyperImpute (jarrett2022hyperimpute) tests on $13$ UC Irvine (UCI) datasets (Kelly2024UCI), a subset of OpenML, with $3$ missingness patterns, and GAIN (gain2018) tests on $6$ UCI datasets with $1$ missingness pattern.
Our code and implementation details for all our contributions above can be accessed on GitHub. [^2] https://anonymous.4open.science/r/tabular-6F65/README.md
Figure 3: Overview of our contributions. The first row demonstrates TabPFN’s imputation method, which performs iterative column-by-column imputation. We build on this by introducing an entry-wise featurization, as shown in the second row. We create a new synthetic data-generator for missingness data to train our model, TabImpute, shown in green (Sec. 3.2 and Sec. 3.3, respectively). We evaluate all the imputers on the comprehensive and rich set of OpenML datasets with many missingness patterns applied (Sec. 4).
@@ 1.1 Previous Work
We build primarily on missing data imputation and tabular representation learning (TRL). While missing data imputation is well-studied with established theory dating back to the 1970s (rubin1976inference), TRL using tabular foundation models is relatively new (muller2022transformers; zhang2025tabpfnmodelruleall; hollmann2023tabpfn; hollmann2025accurate). Below, we describe relevant work that we directly compare against or build upon.
@@
HyperImpute (jarrett2022hyperimpute) combines the power of multiple classical imputation methods through iterative imputation (IM). IM loops over each column, using other columns to predict missing values until convergence. HyperImpute optimizes the imputer at each iteration over candidate methods, including MICE (royston2011multiple), SoftImpute (hastie2015matrix), column mean, MissForest (missforest2011), and Optimal Transport (muzellec2020missing). While achieving great imputation accuracy across diverse missingness patterns and supporting categorical variables, HyperImpute is designed for supervised learning settings and cannot handle matrices with entirely missing columns (common in causal inference and panel data (agarwal2023causal)). Additionally, optimization requires significantly more time than base methods.
@@
Tabular representation learning focuses on building models that can generalize across diverse tabular domains. The pioneering work in this area is TabPFN, a tabular foundation model for supervised learning. TabPFN was proposed in hollmann2023tabpfn and subsequently improved in hollmann2025accurate. Since TabPFN’s introduction, numerous variants have emerged to address scalability and performance limitations, as well as further work aimed at clarifying its internal representations (zhang2025tabpfnmodelruleall; ye2025closer). Recent advances include TabICL (qu2025tabicl), a scalable foundation model that extends supervised learning capabilities to datasets with up to 500K samples through a novel two-stage architecture with column-then-row attention mechanisms. Other notable models include MITRA (MITRA), a tabular foundation model pre-trained on purely synthetic data from a mix of random classifiers/regressors, and Toto (cohen2025timedifferentobservabilityperspective), which is optimized for time series forecasting on observability metrics. Additionally, CausalFM (ma2025foundationmodelscausalinference) enables Bayesian causal inference through structural causal model priors similar to TabPFN, while DO-PFN (robertson2025dopfnincontextlearningcausal) extends the PFN framework to estimate counterfactual distributions from observational data.
@@ 2 Background on Prior-data Fitted Networks and TabPFN
Posterior predictive modeling and synthetic prior fitting.
PFNs are rooted in Bayesian supervised learning, where the primary objective is to model the posterior predictive distribution (PPD) (mackay1992practical; seeger2004gaussian; blei2017variational). Since computing the PPD is often intractable (mackay1992practical), PFNs instead learn to approximate the PPD offline through a process called synthetic prior fitting muller2022transformers. This is achieved using a prior specified by a sampling scheme that first samples a data-generating mechanism, $\phi\sim p(\phi)$, and then samples a synthetic dataset, $D\sim p(D|\phi)$. This process is repeated to generate millions of diverse datasets for training. The network’s parameters, $\theta$, are then optimized to predict held-out test samples ($D_{\text{test}}\subset D$) conditioned on the rest of the dataset ($D_{\text{train}}=D\setminus D_{\text{test}}$). The training objective is to minimize the negative log likelihood (NLL) loss on these held-out examples:
$\mathcal{L}_{\text{NLL}}(\theta)=\mathbb{E}_{((x_{\text{test}},y_{\text{test}})\cup D_{\text{train}})\sim p(D)}[-\log q_{\theta}(y_{\text{test}}|x_{\text{test}},D_{\text{train}})].$ (2)
This training process makes it explicit that PFNs are trained to emulate Bayesian inference by averaging over tasks drawn from a prior distribution over data-generating functions. Minimizing this loss ensures that the trained neural network, $q_{\theta}$, learns to approximate the true Bayesian PPD for the specified prior.
@@
TabPFN is a PFN built specifically for tabular supervised learning (hollmann2023tabpfn). The model employs a novel two-way attention mechanism specifically designed for tabular data. Unlike standard transformers (vaswani2023attentionneed) that treat tabular data as sequential tokens, TabPFN assigns separate representations to each cell in the table. The architecture uses alternating attention patterns: each cell first attends to other features within its row (inter-feature attention), then attends to the same feature across all rows (inter-sample attention). This design ensures permutation invariance for both samples and features while enabling efficient scaling to larger tables than those seen during training. TabPFN v2 (hollmann2025accurate) retains the core training paradigm of the original TabPFN while introducing several key enhancements that improve accuracy, runtime, and applicability. Going forward, when referring to TabPFN, we mean TabPFN v2 since it is the most up-to-date version.
3 Training TabImpute on Synthetic Data
@@
We develop a featurization for tabular missing data imputation that enables better utilization of TabPFN’s architecture, create a synthetic data generation pipeline across diverse missingness patterns, and employ an adaptive training algorithm to balance performance across all patterns. Training used $8$ H200 GPUs and an Intel Xeon Platinum 8592+ CPU over approximately 2 days, processing $1.9$ million synthetic tables. Our model matches TabPFN’s size and runs on CPU-only systems. Evaluation used $1$ H200 GPU.
3.1 Entry-wise Featurization and Architecture
We recast missing data imputation as supervised learning to leverage TabPFN’s architecture and enable parallel GPU computation of missing entries. For each dataset (training point), let $X^{*}$ be the complete matrix with $m$ rows and $n$ columns, $\Omega$ be the set of missing entry indices, and $X$ be the matrix with induced missingness:
$X_{ij}=\begin{cases}X^{*}_{ij}&\text{for}\quad(i,j)\in[m]\times[n]\setminus\Omega\\ \star&\text{otherwise.}\end{cases}$ (3)
@@ 3.1 Entry-wise Featurization and Architecture
where $\star$ denotes a missing entry. Let $\Omega_{\text{obs}}=[m]\times[n]\setminus\Omega$. Our feature matrix construction adds $(i\oplus j\oplus X_{i,:}\oplus X_{:,j})$ for each entry $i,j\in[m]\times[n]$, where $X_{i,:}$ denotes the $i$-th row, $X_{:,j}$ the $j$-th column, and $\oplus$ concatenation. Each row’s target is $y_{ij}=X^{*}_{ij}$. During pre-training, we train the model to predict target values for all $(i,j)\in\Omega$. This procedure creates a feature matrix of size $nm\times(n+m)$. This featurization captures all necessary information for each cell through its row and column context while enabling parallel computation of missing entries on GPUs. Although the input matrix size increases, parallelization gains outweigh this cost.
@@
We use TabPFN’s base architecture with one modification: removing the attention mask to allow input rows to attend to query rows. Since our input/query row sizes vary randomly with missingness patterns (unlike TabPFN’s controlled synthetic generation), we remove the mask to enable parallel batch training. TabPFN’s mask prevents train points from seeing test feature distributions, which is important in general supervised learning. However, our test set is created using data already available to the observed points, thus alleviating any data-leakage concerns.

### v4

# TabImpute: Universal Zero-Shot Imputation for Tabular Data

@@ Abstract
Missing data is a widespread problem in tabular settings. Existing solutions range from simple averaging to complex generative adversarial networks, but due to each method’s large variance in performance across real-world domains and time-consuming hyperparameter tuning, no universal imputation method exists. This performance variance is particularly pronounced in small datasets, where the models have the least amount of information. Building on TabPFN, a recent tabular foundation model for supervised learning, we propose TabImpute, a pre-trained transformer that delivers accurate and fast zero-shot imputations, requiring no fitting or hyperparameter tuning at inference time. To train and evaluate TabImpute, we introduce (i) an entry-wise featurization for tabular settings, enabling a $100\times$ speedup over the previous TabPFN imputation method, (ii) a synthetic training data generation pipeline incorporating a diverse set of missingness patterns to enhance accuracy on real-world missing data problems, and (iii) MissBench, a comprehensive benchmark with $42$ OpenML tables and $13$ new missingness patterns. MissBench spans domains such as medicine, finance, and engineering, showcasing TabImpute’s robust performance compared to numerous established imputation methods.
Machine Learning, Imputation, Matrix Completion, Transformer
@@ 1 Introduction
Missing data is ubiquitous across tabular datasets, affecting statisticians, economists, health officials, and businesses. For example, healthcare datasets may lack some rerecorded blood pressure measurements, or datasets merged from multiple sources may share only features. Regardless of the source, missing data must usually be imputed to numerical values before employing statistical or machine learning models. Imputation methods range from simple approaches (e.g., constant values and averages) to more sophisticated techniques such as nearest-neighbor methods (Batista01052003), matrix factorization methods such as SoftImpute (hastie2015matrix), and random forest regression, notably the MissForest algorithm (missforest2011). However, each method in the literature is typically tailored for specific settings, with performance varying significantly across datasets, domains, and missingness patterns (van2012flexible; jarrett2022hyperimpute; agarwal2023causal; Ibrahim01032005). Building on recent advances in tabular representation learning (hollmann2023tabpfn; ye2025closer), we propose a pre-trained transformer specifically designed for the tabular missing-data problem that produces accurate and fast zero-shot imputations.
@@
Figure 1: Evaluation on real-world OpenML small datasets: MissBench. We compare TabImpute with $13$ other popular methods on MissBench. In panel (a), we plot the imputation accuracy (defined as 1 - normalized RMSE), which is calculated for each method, normalized within a dataset, and averaged across datasets and $13$ missingness patterns. Accuracy scores are normalized to 0-1, with higher values better. Error bars indicate 95% confidence intervals. In panel (b), we compare the runtime per table entry. Any method not labeled (GPU) is tested on a CPU because it is not implemented on GPUs. See Sec. 3 for our exact computing specifications and Sec. 4 for accuracy score methodology.
@@
(rubin1976inference) proposed analyzing missingness based on its relationship with the data-generating process to determine whether missingness biases downstream estimation. Rubin demonstrated that when missingness is independent of the underlying data, the observed data distribution provides an unbiased basis for estimation. This framework categorizes missingness into three classes: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR) (van2012flexible; sportisse2020aimputation; sportisse2020b). MCAR defines scenarios where missingness occurs uniformly and independently of all data values. MAR encompasses cases where missingness depends on observed variables that can be appropriately conditioned upon during analysis. MNAR describes situations where missingness depends on unobserved factors that cannot be easily conditioned on.
@@
While Rubin’s framework characterizes how missingness relates to the underlying data-generating process, it largely treats the missingness indicators for each variable as conditionally independent across entries. Recent work on structured missingness argues that, especially in large-scale multi-source datasets, missingness itself can exhibit rich multivariate structure that is not captured by the standard MCAR/MAR/MNAR taxonomy. In particular, (mitra2023learning) introduces structured missingness as an umbrella term for mechanisms in which missingness follows systematic patterns. (jackson2023completecharacterisationstructuredmissingness) provides a complementary characterization that accounts for dependencies among missingness indicators across variables. We incorporate structured missingness by testing on 11 MNAR patterns.
@@
Previous work typically introduced new missingness patterns within these categories and proposed pattern-specific solutions. For instance, (agarwal2023causal) proposed a nearest neighbor-based matrix completion method specifically designed for a block-wise MNAR pattern. We instead develop a single method to perform well across diverse patterns and data domains by building on TabPFN, a popular tabular foundation model for supervised learning.
@@
TabPFN is a pre-trained transformer model for supervised learning that performs well without any fine-tuning (hollmann2025accurate). The team behind TabPFN created an imputation method in their tabpfn-extensions Python package by using the TabPFN model in iterative column-wise imputation, denoted here as Col-TabPFN (available on GitHub [^1] https://github.com/PriorLabs/tabpfn-extensions/blob/main/src/tabpfn_extensions/unsupervised/unsupervised.py). However, when evaluated on our new benchmark MissBench, consisting of $42$ real-world small OpenML datasets and $13$ missingness patterns, Col-TabPFN’s imputation approach shows low accuracy and slow runtime. We improve on this approach by introducing a new entry-wise featurization (EWF), allowing parallel prediction of each missing value using TabPFN’s model, denoted EWF-TabPFN. To improve on EWF-TabPFN, we train a new underlying model specifically designed for tabular data imputation, TabImpute, to better fit this class of tasks, achieving the results shown in Fig. 1 and Tab. 1.
@@
  • We propose a new state-of-the-art pre-trained transformer model, TabImpute, for missing data imputation in small tables by building on TabPFN and introducing a new entry-wise missing data featurization (see Sec. 3 for details).
  • We develop a synthetic data generation pipeline to create training datasets with a comprehensive collection of missing values covering a wide range of realistic MNAR patterns (see Sec. 3.1, Sec. A.2, and Tab. 10 for details).
  • We demonstrate TabImpute’s performance on a novel, comprehensive test bench of small tables, which we call MissBench, using $42$ real-world small OpenML datasets and $13$ missingness patterns (see Sec. 4 for details) spanning many fields including biology, medicine, finance, and education.
Our code and implementation details for all our contributions above can be accessed on GitHub. [^2] https://github.com/jacobf18/tabular
Figure 3: Overview of our contributions. The first row demonstrates Col-TabPFN, which performs iterative column-by-column imputation. We build on this by introducing an entry-wise featurization, as shown in the second row. We create a new synthetic data-generator for missingness data to train our model, TabImpute, shown in green (Sec. 3.1 and Sec. 3.2, respectively). We evaluate all the imputers on the comprehensive and rich set of OpenML datasets with many missingness patterns applied (Sec. 4).
@@ 1.1 Previous Work
We primarily build on missing-data imputation and tabular representation learning (TRL). While missing-data imputation is well-studied, with established theory dating back to the 1970s (rubin1976inference), TRL using tabular foundation models is relatively new (muller2022transformers; zhang2025tabpfnmodelruleall; hollmann2023tabpfn; hollmann2025accurate). Below, we describe relevant work.
@@
HyperImpute (jarrett2022hyperimpute) combines the power of multiple classical imputation methods through iterative imputation (IM). IM loops over each column, using other columns to predict missing values until convergence. HyperImpute optimizes the imputer at each iteration over candidate methods, including MICE (royston2011multiple), SoftImpute (hastie2015matrix), column mean, MissForest (missforest2011), and Optimal Transport (muzellec2020missing).
@@
Tabular representation learning focuses on building models that can generalize across diverse tabular domains. The pioneering work in this area is TabPFN, proposed in (hollmann2023tabpfn) and improved in (hollmann2025accurate). Since TabPFN’s introduction, numerous variants have emerged to address scalability and performance limitations, as well as further work aimed at clarifying its internal representations (zhang2025tabpfnmodelruleall; ye2025closer). Recent advances include TabICL (qu2025tabicl), a scalable foundation model that extends supervised learning capabilities to datasets with up to 500K samples through a novel two-stage architecture with column-then-row attention mechanisms. Other notable models include MITRA (MITRA), a tabular foundation model pre-trained on purely synthetic data from a mix of random classifiers/regressors, and Toto (cohen2025timedifferentobservabilityperspective), which is optimized for time series forecasting on observability metrics. Additionally, CausalFM (ma2025foundationmodelscausalinference) enables Bayesian causal inference through structural causal model priors similar to TabPFN, while DO-PFN (robertson2025dopfnincontextlearningcausal) extends the PFN framework to estimate counterfactual distributions from observational data.
@@ 2 Background on Prior-data Fitted Networks and TabPFN
Posterior predictive modeling.
PFNs are rooted in Bayesian supervised learning, where the primary objective is to model the posterior predictive distribution (PPD) (mackay1992practical; seeger2004gaussian; blei2017variational). Since computing the PPD is often intractable (mackay1992practical), PFNs instead learn to approximate the PPD offline through a process called synthetic prior fitting (muller2022transformers). This is achieved using a prior specified by a sampling scheme that first samples a data-generating mechanism, $\phi\sim p(\phi)$, and then samples a synthetic dataset, $D\sim p(D|\phi)$. This process is repeated to generate millions of diverse datasets for training. The network’s parameters, $\theta$, are then optimized to predict held-out test samples ($D_{\text{test}}\subset D$) conditioned on the rest of the dataset ($D_{\text{train}}=D\setminus D_{\text{test}}$). The training objective is to minimize the negative log likelihood (NLL) loss on these held-out examples: $\mathcal{L}(\theta)=\mathbb{E}_{((x_{\text{test}},y_{\text{test}})\cup D_{\text{train}})\sim p(D)}[-\log q_{\theta}(y_{\text{test}}|x_{\text{test}},D_{\text{train}})].$ Minimizing this loss ensures that the trained neural network $q_{\theta}$ approximates the true Bayesian PPD for the specified prior.
@@
TabPFN is a PFN built specifically for tabular supervised learning (hollmann2023tabpfn). The model employs a novel two-way attention mechanism specifically designed for tabular data. The architecture uses alternating attention patterns: each cell first attends to other features within its row (inter-feature attention), then attends to the same feature across all rows (inter-sample attention). This design ensures permutation invariance for both samples and features while enabling efficient scaling to larger tables than those seen during training. TabPFN v2 (hollmann2025accurate) retains the core training paradigm of TabPFN v1 while introducing several key enhancements that improve accuracy, runtime, and applicability. Going forward, when referring to TabPFN, we mean TabPFN v2.
3 Pre-Training TabImpute on Synthetic Data
@@
We pre-train the model on our novel missing-data prior using entry-wise featurization, both detailed below. Training used $8$ H200 GPUs and an Intel Xeon Platinum 8592+ CPU over approximately 30 hours, processing $1.9$ million synthetic tables, and evaluation used $1$ H200 GPU. Our model matches TabPFN’s size and runs on CPU-only systems.
Entry-wise Featurization.
@@
Let $X^{*}$ be the complete matrix with $m$ rows and $n$ columns, $\Omega$ be the set of missing entry indices, and $X$ be the matrix with induced missingness: for each entry, $X_{ij}=X^{*}_{ij}$ if entry $(i,j)\notin\Omega$ (observed) and $X_{ij}=\star$ if $(i,j)\in\Omega$ (missing) where $\star$ denotes a missing entry. Let $\Omega_{\text{obs}}=[m]\times[n]\setminus\Omega$. Each row of the entry-wise feature matrix is $(i\oplus j\oplus X_{i,:}\oplus X_{:,j})$ for entry $i,j\in[m]\times[n]$, where $X_{i,:}$ denotes the $i$-th row, $X_{:,j}$ the $j$-th column, and $\oplus$ concatenation. Each row’s target is $y_{ij}=X^{*}_{ij}$. During pre-training, we train the model to predict target values for all $(i,j)\in\Omega$. This procedure creates a feature matrix of order $nm\times(n+m)$. This featurization captures all necessary information for each cell through its row and column context while enabling parallel computation of missing entries on GPUs. Although the input matrix size increases, the gains from parallelization outweigh this cost for small tables. Note that the features are created using the matrix with missing values $X$, not the true matrix $X^{*}$.
We demonstrate entry-wise featurization on a small toy matrix with missing values in Fig. 4. During pre-training, we have access to the true values of the $\mathrm{NaN}$s in $y$. However, these values are only used to update the model’s weights via the loss function; the true values are never input into the model. The model internally handles $\mathrm{NaN}$ inputs by replacing their values with the column mean in the input matrix and then appending a binary matrix indicating which values are $\mathrm{NaN}$. Internally, the $\mathrm{NaN}$ values in $y$ are replaced with the mean of the non-$\mathrm{NaN}$ values (observed values) in $y$.
[…]
Figure 4: Entry-wise featurization to create a dataset pair $(X_{\text{EWF}},y)$: each entry of the original matrix $X$ produces one row in the feature matrix $X_{\text{EWF}}$, with the corresponding target $y_{(r,c)}=X_{r,c}$.
@@
Following TabPFN’s terminology, we refer to rows in the feature matrix with observed target values (non-$\mathrm{NaN}$) as input points and rows with missing values as query points. We adopt TabPFN’s base transformer architecture with one modification: we remove the attention mask that restricts information flow between query points. In the original supervised learning setting considered by TabPFN, this mask encodes an inductive bias that predictions for query points should be conditionally independent, reflecting the standard assumption that test points do not influence each other’s predictions. In contrast, in missing-data imputation, there is no assumption that missing entries (query points) should be independent. Indeed, under many missingness mechanisms, particularly non-MCAR settings, missing values may be strongly correlated through shared latent structure. Allowing query points to exchange information, therefore, better reflects the structure of the imputation problem.